OpenStudy (anonymous):

prove sinh2x=2 sinhx.coshx

OpenStudy (anonymous):

man I already proved this.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@adam160596 did you forget your other post?

OpenStudy (anonymous):

$$\sinh x=\frac12(e^x-e^{-x})\\\cosh x=\frac12(e^x+e^{-x})\\\sin2x=\frac12(e^{2x}-e^{-2x})=\frac12\left((e^x)^2-(e^{-x})^2\right)=\frac12(e^x+e^{-x})(e^{x}-e^{-x})=2\sinh x\cosh x$$

OpenStudy (anonymous):

sorry i put in the wrong question :/

OpenStudy (jhannybean):

xD